# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v2

@@ 6 Experiments
    Intensity-free approaches. We implement IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024).
  3. 3.
@@
    Intensity-based approaches. We implement RMTPP (Du et al., 2016) as an example of the TPP approach with a traditional RNN. We add NHP (Mei & Eisner, 2017), based on a continuous time LSTM architecture. We also evaluate an AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019).
  4. 4.
    Next-K approaches. We adapt IFTPP and RMTPP to predict multiple future events directly, implementing IFTPP-K and RMTPP-K, respectively. These approaches originate in time series analysis and have not been previously applied in the MTPP domain.
  5. 5.
@@
    Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses and selects the best sequence using a contrastive approach.
@@
Additional details on these methods are provided in Appendix A, with training specifics outlined in Appendix D. Hyperparameters for both the methods and metrics are presented in Appendix E. The main evaluation results are shown in Table 1. In the following sections, we will discuss these results from various perspectives and offer additional analysis of the methods’ behavior.
6.1 Metrics Comparison
@@
The evaluation results in Table 1 show that the OTD metric can yield low values even for simple rule-based baselines. For instance, the Last baseline achieves the lowest OTD values on the Retweet and Amazon datasets. In contrast, the T-mAP metric more clearly distinguishes between rule-based baselines and deep learning models. This difference arises from the ability to measure model confidence. OTD evaluates predicted labels alone, while T-mAP measures average precision by adjusting logits thresholds. Unlike rule-based baselines, deep models provide confidence scores, whereas baselines output only labels, resulting in low T-mAP scores for simple baselines.
The HYPRO and NHP methods maximize T-mAP, while the lowest OTD values are achieved by statistical baselines, IFTPP, and RMTPP. The difference between the two metrics lies in balancing the importance of event order versus time prediction quality. OTD evaluates the first 5-10 events, and to minimize OTD, a model must correctly identify the events that constitute the beginning of the sequence. IFTPP and RMTPP jointly predict the next event’s time and label, leading to better ordering. In contrast, methods like NHP, AttNHP, and ODE predict time steps independently for each event type. When the differences between timestamps are small, these methods can result in random ordering, negatively affecting the OTD score. T-mAP, on the other hand, is less sensitive to event order as long as the time predictions are accurate.
@@
The HYPRO method, designed to differentiate between ground truth and predicted sequences using a discriminator model, achieves high T-mAP scores. This highlights T-mAP’s ability to reward models that generate realistic sequences. We also observed that methods with high T-mAP scores often have low next-event prediction accuracy. For instance, on the Retweet, Amazon, and MIMIC-IV datasets, high long-horizon performance is often accompanied by low next-event prediction scores. This is particularly true for the HYPRO method, whose loss function is heavily focused on long-horizon quality. The contrast on the MIMIC-IV dataset can be attributed to the large number of events with identical timestamps, making precise event ordering more difficult and impacting both next-event and OTD scores.
6.2 Intensity-Based vs Intensity-Free
Previous works mostly compared either intensity-based neural methods (Du et al., 2016; Xue et al., 2022) or intensity-free approaches (Padhi et al., 2021; McDermott et al., 2024). The only exception is IFTPP (Shchur et al., 2019), which did not compare with NHP and included only test set likelihood as an evaluation metric. This raises the question: which approach is superior? We compare the intensity-free IFTPP method with intensity-based RMTPP, NHP, and ODE approaches. Table 1 shows that the intensity-free IFTPP method excels in the next-event MAE prediction, as it optimizes this metric. In other scenarios, the results vary between datasets. Therefore, we conclude that there is no preferred solution, and both approaches warrant attention in future research.
6.3 Autoregression vs Direct Horizon Prediction
We compared the IFTPP and RMTPP methods with their Next-K counterparts. Our results indicate that autoregressive models generally perform better regarding next-event prediction quality, while Next-K approaches improve T-mAP scores on all datasets except Transactions. The OTD-based rankings vary across datasets, with no single approach consistently outperforming the others. Despite these findings, little effort has been made to adapt Next-K models to MTPPs, suggesting future research in this area.
We also observed that the Next-K approaches exhibit some interesting properties compared to autoregression. For instance, the entropy of label distributions in autoregression models decreases with increasing generation steps, as shown in Figure 5. The likely reason for this behavior is the dependency of the model’s future output on its past errors. Conversely, Next-K models demonstrate increasing entropy, which is expected as future events become harder to predict. This suggests that Next-K models have the potential to predict better the distributions of future labels, a factor that should be considered in future research.
Figure 5: Entropy of label distributions as a function of the position in the generated sequence. Results for StackOverflow and Amazon datasets are provided in Appendix F.
6.4 The Maximum Sequence Length
We observed that long-horizon prediction quality largely depends on the maximum number of allowed predictions in autoregression and Next-K models. As shown in Fig. 6, the optimal number of predicted events is usually less than the maximum number of events on the horizon. For example, in the Transactions dataset, the optimal length ranges from 3 to 5, depending on the method. This indicates that the quality of confidence estimation degrades for events further in the future, therefore it is beneficial to limit the number of predictions manually.
Figure 6: T-mAP dependency on the maximum length of the predicted sequence. Results for the StackOverflow and Amazon datasets are detailed in Appendix G.
Therefore, we conclude that greater attention should be given to probability calibration between generation steps. This finding also highlights progress in long-horizon prediction tasks, particularly determining the maximum horizon that can be accurately modeled. T-mAP is the only metric capable of evaluating the optimal predicted sequence length, while the OTD metric remains unaffected by sequence length as it compares fixed-size prefixes.
7 Limitations and Future Work
@@
The proposed benchmark has several shortcomings. The training process for some methods is relatively slow. For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future. Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training. Additionally, the list of implemented backbone architectures and losses can be extended, for example, by incorporating hybrid (Deshpande et al., 2021) and diffusion models (Zhou et al., 2023).
Our benchmark encourages future research to develop improved techniques for predicting future events and establish simple baselines for better measuring progress in the field. For example, Next-K models, which predict multiple future events without autoregression, show promising results in long-horizon prediction tasks. We suggest further exploration of these models. We also highlight the importance of label distribution estimation and emphasize the need to improve confidence estimation and calibration.
@@
The T-mAP metric can be applied in domains beyond event sequences, like action recognition (Su & Hua, 2017). T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW. T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error. Our theoretical justification of T-mAP can potentially be adapted to the mAP estimation algorithms in computer vision.
8 Conclusion
@@
In this paper, we propose the HoTPP benchmark to assess the quality of long-horizon events forecasting. Our extensive evaluations using established datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction does not necessarily correlate with superior performance in horizon prediction. This finding underscores the need for benchmarks like HoTPP, which emphasize the importance of long-horizon accuracy and robustness. By shifting the focus from short-term to long-term predictive capabilities, HoTPP aims to drive the development of more sophisticated and reliable event sequence prediction models. This, in turn, has the potential to significantly enhance the practical applications of sequential event prediction in various domains, fostering innovation and improving decision-making processes across a wide range of industries.
Appendix A Background on Marked Temporal Point Processes Modeling
@@
$\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$ (5)
@@
Recent research has focused on modeling complex intensity functions. Various neural network architectures have been adapted to address this problem. These approaches differ in the type of neural network used and the model of the intensity function. Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei & Eisner, 2017) and Neural ODEs (Rubanova et al., 2019). The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
Next-K models. While not commonly used in the MTPP field, Next-K approaches are popular in time series modeling (Lim & Zohren, 2021). These methods predict multiple future events at once, avoiding the need for autoregressive inference. The advantages of Next-K approaches include fast inference and stability, as predictions do not depend on potential errors from previous steps, unlike autoregression. The main limitation is a fixed prediction horizon, as the model cannot predict sequences of arbitrary length. However, applying a modified autoregressive approach can potentially address this limitation.
In our work, we implement a Next-K variant of the IFTPP model, which is straightforward. Given K predictions and K ground truth events, we compute MAE and cross-entropy losses between corresponding pairs of events from both sequences. The Next-K variant of the RMTPP approach is more complex, as it violates Hawkes assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1. Consequently, RMTPP-K, unlike RMTPP, cannot model dependencies between the predicted K events. Despite this, RMTPP-K still performs strongly, particularly on the StackOverflow dataset.
@@ Appendix B T-mAP Computation and Proofs
Figure 7: Illustration of the iterative process in Lemma A.
If the size of $m^{\prime}$ is greater than the size of matchings in $M(\mathcal{G})$ then the lemma holds. The optimal matching in the extended graph $\mathcal{G}^{\prime}$ cannot have a size smaller than the optimal matching in $\mathcal{G}$. Now, consider the case when both matchings have the same size. We will show that $m^{\prime}$ has a total weight equal to that of the matchings in $M(\mathcal{G})$.
@@
If the process finishes at a predicted event, then $m^{\prime}$ has a size greater than the size of $m$, and the lemma holds. If the process finishes at a ground truth event, we can replace a part of matching $m^{\prime}$ traced with a part of matching $m$. The resulting matching $\hat{m}$ will have a size equal to both $m$ and $m^{\prime}$ and will have a total weight equal to both matchings. This proves the lemma. ∎
@@
$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (6)
@@
Denote $i$ the only vertex present in $\mathcal{G}$ and filtered from $\mathcal{G}_{h}$. If there is no edge containing vertex $i$ in the matching $m(\mathcal{G})$, then $m(\mathcal{G})$ is also the optimal matching in $\mathcal{G}_{h}$, and theorem holds.
Consider the case when matching $m(\mathcal{G})$ contains an edge $(i,j)$. Let $\hat{m}_{h}=m(\mathcal{G})\setminus\{(i,j)\}$, i.e., the matching without edge $(i,j)$. If it is optimal for the graph $\mathcal{G}_{h}$, then it satisfies the theorem. Otherwise, there is an optimal matching $m_{h}$ with either (a) more vertices or (b) a smaller total weight than $\hat{m}_{h}$.
In (a), matching $m_{h}$ has a size greater than the size of $\hat{m}_{h}$. It follows, that $m_{h}$ has the maximum size in the full graph $\mathcal{G}$. If the total weight of $m_{h}$ is larger or equal to the optimal weight in $\mathcal{G}$, then it can not be less than the total weight of $\hat{m}_{h}$, which contradicts our assumption. At the same time, the total weight of $m_{h}$ can not be less than the optimal weight in the full graph $\mathcal{G}$. It follows that case (a) leads to a contradiction.
@@
$\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$ (7)
@@
$\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$ (8)
@@
We have several correct and incorrect predictions for each threshold $h$ on the predicted label probability. These quantities define precision and recall values along with the total number of ground truth events. A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|\leq\delta$. Note that each target can be assigned to at most one prediction. Therefore, we define a matching, i.e., the correspondence between predictions and ground truth events. The theorem states that the maximum size matching for each threshold $h$ can be constructed as a subset of an optimal matching between full sequences $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and $\mathcal{S}^{i,l}_{gt}$, where optimal matching is a solution of the assignment problem with a bipartite graph, defined in Section 4.2.
@@
$C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$ (9)

### v3

@@ 4.2. Methods
Intensity-free approaches. We implement the IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024). IFTPP uses GRU reccurrent network as a backbone (Cho et al., 2014). We also evaluate IFTPP-T model with a transformer backbone (Radford et al., 2019).
@@
Intensity-based approaches. We implement RMTPP (Du et al., 2016) as an example of the TPP approach with a traditional RNN. We add NHP (Mei and Eisner, 2017), based on a continuous time LSTM architecture. We also evaluate the AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019), which applies ordinary differential equations for temporal dynamics modeling.
@@
Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses with RMTPP and selects the best sequence using a contrastive approach.
Diffusion. We include the Diffusion approach (Zhou et al., 2025), which extends previous research in language modeling (Li et al., 2022). Diffusion model iteratively refines predicted sequence, simultaneously generating all events within the specified horizon. Notably, HoTPP provides the first public implementation of this approach.
We also explored ContiFormer (Chen et al., 2023) and NJDTPP (Zhang et al., 2024), but excluded them from our final comparison for the reasons detailed in Appendix A.1.
Additional details on implemented methods are provided in Appendix A, with training specifics outlined in Appendix D and hyperparameters listed in Appendix E.
5. Experiments
@@
The main evaluation results are presented in Table 2. In the following sections, we will discuss these results from various perspectives and offer additional analysis of the methods’ behavior.
5.1. Long-horizon Prediction Quality
According to Table 2, autoregressive approaches such as IFTPP, RMTPP, and NHP excel at next-event prediction (MAE and accuracy), while horizon prediction methods like HYPRO and Diffusion perform better in terms of OTD and T-mAP. These findings indicate that next-event and horizon prediction are distinct tasks requiring specialized models, and none of the existing approaches achieve top performance in both domains.
Our statistical baselines rank first in 3 out of 10 long-horizon prediction tasks, and in three additional tasks they perform on par with the top methods without a statistically significant difference. These findings highlight the need for greater focus on long-horizon forecasting and point to opportunities for future improvements.
Intensity-based approaches such as RMTPP and NHP generally outperform the autoregressive, intensity-free strategy. Meanwhile, Diffusion relies on intensity-free losses, suggesting that integrating intensity-based loss functions into Diffusion could potentially enhance long-horizon prediction.
Notably, Transformer-based methods rarely offer advantages over RNN architectures. For instance, IFTPP-T significantly surpasses the RNN-based IFTPP only on MIMIC-IV, while being outperformed in most other scenarios. Consequently, the potential of Transformer architectures in the MTPP domain remains insufficiently explored.
Figure 5. Comparison of predicted label entropies, with the dashed horizontal line indicating the ground truth entropy.
5.2. Metrics Comparison
@@
Table 2 shows that the OTD metric can yield low values even for simple rule-based baselines. For example, the MostPopular baseline achieves the lowest OTD scores on the Retweet and Amazon datasets. In contrast, T-mAP more clearly distinguishes between rule-based baselines and deep learning models by measuring predicted labels probabilities rather than hard labels.
The difference between OTD and T-mAP also emerges when comparing autoregressive methods that use the NHP loss (NHP, ODE) with those that jointly model timestamp and label for the next event (IFTPP, RMTPP). For instance, NHP performs poorly on OTD yet ranks among the top autoregressive approaches by T-mAP. This discrepancy arises from how the two metrics balance event order and time prediction quality. OTD considers only the first 5–10 events, so the model must accurately predict these initial events. Because IFTPP and RMTPP jointly predict the next event’s time and label, they generally preserve ordering more effectively. In contrast, NHP, AttNHP, and ODE predict times separately for each event type, causing random ordering when timestamps are close and thereby lowering the OTD score. We conclude that T-mAP is less sensitive to event order as long as the predicted times are accurate.
@@
The HYPRO and Diffusion methods, specifically designed for long-horizon prediction, achieve high T-mAP scores but often underperform autoregressive approaches in terms of OTD. This highlights T-mAP’s ability to reward models that produce realistic sequences. We also observe that methods with high T-mAP scores frequently have low next-event prediction accuracy. For example, on the Retweet, Amazon, and StackOverflow datasets, high long-horizon performance often coincides with low next-event prediction scores. This is especially true for HYPRO and Diffusion, whose loss functions heavily emphasize long-horizon quality. Meanwhile, the strong correlation of next-event prediction quality and long-horizon performance on MIMIC-IV can be attributed to the large number of events with identical timestamps, which makes precise event ordering crucial and favors autoregressive approaches like IFTPP.
Table 3. Examples of predicted sequences (labels only).
Method Transactions MIMIC-IV Retweet Amazon StackOverflow IFTPP 6, 6, 6, 6, 6, 6 14, 12, 4, 6, 11, 10, 1, 3, 5, 8, 9, 16 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 11, 0, 11, 0, 0, 0, 0, 0, 0, 0, 0 7, 17, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3 RMTPP 6, 6, 6, 6, 6, 6 14, 4, 6, 11 0, 0, 0, 0, 0, 0, 0 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11 17, 3, 3, 3, 3, 3 NHP 6, 6, 6, 6, 6 4, 6 1, 1, 1, 1, 1, 1, 1, 1 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 AttNHP 1, 1, 1, 1, 1 4, 6, 10, 1 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 3, 0, 4, 1, 1, 1, 0, 0, 0, 0, 0, 0 3, 3, 3, 3, 3, 3, 3 ODE 1, 6, 6, 6, 6, 6 4, 6, 1, 3, 5, 2, 7, 4, 6 0, 0, 0, 0, 0, 0, 0 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 HYPRO 3, 6, 6, 1, 6, 1, 6 24, 1, 3, 5, 17, 8, 2, 20 0, 0, 1, 0, 0, 1, 0 1, 11, 0, 0, 11, 11, 11, 5, 1, 0, 5, 11 17, 5, 3, 5, 13, 3, 5 Diffusion 6, 6, 6, 6, 6, 6, 6 14, 4, 6, 1, 1, 3, 3, 3, 8, 3, 2, 2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 0, 0, 8, 5, 11, 11, 5, 5, 0, 5, 5, 11 3, 3, 3, 3, 3, 3, 3, 3
5.3. Mode Collapse
We found that most models tend to predict a single label or repetitive patterns in long-horizon tasks, as shown in Table 3. To quantify this effect, we measured the entropies of predicted labels (Fig.5). Autoregressive methods generally exhibit low prediction diversity across all datasets except MIMIC-IV, where strong dependencies enable the model to learn typical sequences. Meanwhile, HYPRO, which samples from RMTPP with an increased temperature, substantially enhances label diversity. Diffusion similarly boosts diversity except on the Transactions dataset. Overall, most methods, except HYPRO, experience mode collapse in long-horizon prediction, although horizon-based approaches partly mitigate this issue. Sampling with an increased temperature can potentially solve this issue by offering a trade-off between predictions quality and diversity. Additional examples of predicted sequences are presented in Appendix G.
6. Limitations and Future Work
@@
The proposed benchmark has several shortcomings. The training process for some methods is relatively slow. For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future. Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training. Additionally, the list of implemented methods can be extended, for example, by incorporating multi-task models (Deshpande et al., 2021).
Our benchmark encourages future research to develop methods capable of solving next-event and long-horizon prediction tasks simultaneously, as discussed in Section 5.1. One promising direction involves adapting concepts from our statistical baselines for deep models, thereby surpassing baseline performance. A key question in this effort is the correct alignment between predicted and ground truth events during training, as discussed in the context of T-mAP (Section 3). Future work could also target the reduction of mode collapse and the enhancement of predictions diversity, as highlighted in Section 5.3.
@@
On the other hand, the T-mAP metric can be applied in domains beyond event sequences, such as action recognition (Su and Hua, 2017). T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW. T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error. Our theoretical justification of T-mAP can potentially be adapted for mAP estimation algorithms in computer vision.
7. Conclusion
@@
In this paper, we propose the HoTPP benchmark to assess the quality of long-horizon events forecasting. HoTPP provides optimized implementations of the state-of-the-art methods ranging from statistical baselines to ODE and Diffusion (Sec.4), along with diverse metrics for both next-event and long-horizon prediction (Sec.3). Our extensive evaluations using diverse datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction typically does not correlate with superior performance in horizon prediction (Sec. 5.1). Moreover, most existing methods suffer from mode collapse when forecasting longer sequences (Sec. 5.3). These findings underscore the need for developing specialized models. By emphasizing long-term predictive capabilities, HoTPP aims to stimulate the development of more robust and reliable event-sequence models. This advance has the potential to substantially enhance practical applications in various industries, driving both innovation and more informed decision-making.
Appendix A Modeling Marked Temporal Point Processes
@@
(5) $\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$
@@
Recent research has focused on modeling complex intensity functions. Various neural network architectures have been adapted to address this problem. These approaches differ in the type of neural network used and the model of the intensity function. Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei and Eisner, 2017) and Neural ODEs (Rubanova et al., 2019). The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
Diffusion. Recently, diffusion-based approaches have been applied to NLP tasks (Li et al., 2022). This idea has been extended to event sequences by adapting best practices from NLP (Zhou et al., 2025). In such methods, events are first mapped to continuous embeddings, where the diffusion process operates, then mapped back to event space. The process begins with embeddings sampled from Gaussian noise and iteratively refines them using historical event embeddings. Finally, the refined embeddings are projected to event times and labels through a softmax-based layer. Training involves a combination of MSE loss in the continuous space, MAE loss for time reconstruction, and cross-entropy loss for label prediction. In HoTPP, we provide the first public implementation of the approach proposed by (Zhou et al., 2025).
MostPopular Baseline. This baseline estimates the average inter-event time and the empirical label distribution from historical data. For next-event prediction, it outputs the average inter-event duration and selects the label with the highest probability. For long-horizon forecasting, it generates events at fixed time intervals and assigns labels such that the predicted distribution aligns with the estimated historical distribution, thus minimizing the discrepancy between predicted and observed label frequencies. Because it assigns hard labels, this method is particularly designed for next-event prediction and minimizing the OTD metric.
HistoryDensity Baseline. Similar to MostPopular, this approach estimates label distributions from historical data. Unlike MostPopular, it focuses on label densities with respect to time rather than normalizing over the total number of events. Consequently, for long-horizon forecasting, it assigns soft probabilities to each predicted event, thereby maximizing the T-mAP score.
A.1. Methods Excluded from Comparison
We attempted to incorporate ContiFormer (Chen et al., 2023) into our benchmark but encountered multiple challenges. The official implementation provides scripts only for training on a toy spiral dataset, making it impossible to reproduce the method’s performance on real-world data. Moreover, training even on this toy example required approximately three hours on an NVIDIA RTX 4060, suggesting that ContiFormer cannot be efficiently scaled to large datasets such as Transactions. In addition, contrary to the claims in the paper, ODE (Rubanova et al., 2019) and continuous-time transformers (Yang et al., 2022) were already discussed in earlier work. Therefore, we excluded ContiFormer from our benchmark in favor of ODE and AttNHP.
We also attempted to include NJDTPP (Zhang et al., 2024) in our benchmark, but encountered several challenges. Even on StackOverflow, our smallest dataset, the inference time spanned two hours for a single test set prediction. Moreover, after running the official implementation on StackOverflow, we observed accuracy values that were inconsistent with the paper and fell below most baseline results. We also identified data leakage during inference, where future event timestamps were used to predict the subsequent event. Consequently, we chose not to include NJDTPP in our comparisons.
@@ Appendix B T-mAP Computation and Proofs
Figure 6. Illustration of the iterative process in Lemma A.
If the size of $m^{\prime}$ is greater than the size of matchings in $M(\mathcal{G})$, then the lemma holds. The optimal matching in the extended graph $\mathcal{G}^{\prime}$ cannot have a size smaller than the optimal matching in $\mathcal{G}$. Now, consider the case when both matchings have the same size. We will show that $m^{\prime}$ has a total weight equal to that of the matchings in $M(\mathcal{G})$.
@@
If the process finishes at a predicted event, then $m^{\prime}$ has a size greater than the size of $m$, and the lemma holds. If the process finishes at a ground truth event, we can replace a part of matching $m^{\prime}$ traced with a part of matching $m$. The resulting matching $\hat{m}$ will have a size equal to $m$ and $m^{\prime}$ and a total weight equal to both matchings. This proves the lemma. ∎
@@
(6) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$
@@
Denote $i$ to be the only vertex present in $\mathcal{G}$ and filtered from $\mathcal{G}_{h}$. If there is no edge containing vertex $i$ in the matching $m(\mathcal{G})$, then $m(\mathcal{G})$ is also the optimal matching in $\mathcal{G}_{h}$ and theorem holds.
Consider the case when matching $m(\mathcal{G})$ contains an edge $(i,j)$. Let $\hat{m}_{h}=m(\mathcal{G})\setminus\{(i,j)\}$, i.e., the matching without edge $(i,j)$. If it is optimal for the graph $\mathcal{G}_{h}$, then it will satisfy the theorem. Otherwise, there is an optimal matching $m_{h}$ with either (a) more vertices or (b) a smaller total weight than $\hat{m}_{h}$.
In (a), matching $m_{h}$ has a size greater than the size of $\hat{m}_{h}$. It follows that $m_{h}$ has the maximum size in the complete graph $\mathcal{G}$. If the total weight of $m_{h}$ is larger or equal to the optimal weight in $\mathcal{G}$, then it can not be less than the total weight of $\hat{m}_{h}$, which contradicts our assumption. At the same time, the total weight of $m_{h}$ can not be less than the optimal weight in the complete graph $\mathcal{G}$. It follows that case (a) leads to a contradiction.
@@
(7) $\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$
@@
(8) $\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$
@@
We have several correct and incorrect predictions for each threshold $h$ on the predicted label probability. These quantities define precision, recall, and the total number of ground truth events. A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|\leq\delta$. Note that each target can be assigned to at most one prediction. Therefore, we define a matching, i.e., the correspondence between predictions and ground truth events. The theorem states that the maximum size matching for each threshold $h$ can be constructed as a subset of an optimal matching between full sequences $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and $\mathcal{S}^{i,l}_{gt}$, where optimal matching is a solution of the assignment problem with a bipartite graph, defined in Section 3.2.2.
@@
(9) $C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$